One- and two-sided tests We know the male students height is approximately normal, and has standard deviation 4 inches. In a sample of 10 male students, the mean height is 68 inches. Calculate the p value corresponding to the following null hypotheses. Ho The average height of male students in this college is 70 inches. H1 : The average height of male students in this college is not 70 inches. Ho :The average height of male students in this college is at least 70 inches. H1 : The average height of male students in this college less than 70 inches
Answer #1

Ans:

1)

Two tailed test:

Test statistic:

z=(68-70)/(4/sqrt(10))

z=-1.58

p-value(2 tailed)=2*P(z<-1.58)=0.1141

(if exact z value is taken,then p-value=0.1138)

2)Left tailed

p-value(left tailed)=P(z<-1.58)=0.0571

(if exact z value is taken,then p-value=0.0569)